# Colimits of Heyting Algebras through Esakia Duality

### v1

@@ 4.3 Pushouts and Co-Amalgamation
are surjective, since for each $x\in X$, $f(x)=g(y)$ for some $y\in Y$, by surjectivity of $g$, and vice-versa. Certainly then the unique lifting from $V_{G}(X\times_{Z}Y)$ to the projections will likewise be surjective, and both will be p-morphisms, showing that amalgamation holds. ∎
@@ 5 Subvarieties of Heyting algebras
The construction just outlined can also be adapted to handle other varieties of Heyting algebras. In this section we illustrate this by providing analogous free constructions of $\mathsf{KC}$-algebras and $\mathsf{LC}$-algebras; we leave a detailed analysis of the full scope of the method for future work.
@@
they are also sometimes called WLEM-algebras (For Weak Excluded Middle), DeMorgan Algebras, amongst several other names. They appear importantly in settings like Topos Theory (see e.g. [10]). In turn, $\mathsf{LC}$-algebras are Heyting algebras axiomatised by
@@
they are sometimes called “Gödel algebras” or “Gödel-Dummett algebras” as well as “prelinear Heyting algebras” (the latter is due to the fact that the subdirectly irreducible elements in such a variety are chains).
@@ 5.1 KC-algebras
Let $(X,\leq)$ be a Priestley space. Let $*R=(\leq\cup\leq^{\frown})^{*}$.We say that a subset $C$ is order-connected if for each $x,y\in C$, we have $xRy$. We say that $C$ is an order-component if it is a maximal order-connected subset.
@@
Let us first outline our construction in algebraic terms. Given an appropriate distributive lattice $\mathcal{D}$, consider the complex:
@@
where we denote, as usual, the inclusions as $i_{i,j}$. In order to generate a $\mathsf{KC}$ algebra, we need to quotient this complex at appropriate steps. Namely, we need to replace $H^{2}(\mathcal{D})$ by the following:
@@
We then construct the alternative complex:
@@
Note that the limit of this construction will certainly be a Heyting algebra, and indeed, it will be a $\mathsf{KC}$-algebra: for any choice $a$ of value for $p$, there is a step where $\neg\neg a\vee\neg a$ is made equivalent to $\top$. Hence the question lies whether such a construction is still universal.
For that purpose we will find a dual meaning to forming the above quotient.
@@
Let $g:X\to Y$ be a continuous and order-preserving map. Let $C\in V_{r_{g}}(V_{g}(X))$ be an element of the second stage of the $g$-Vietoris complex over $X$. We say that $C$ is well-directed if whenever $D,D^{\prime}\in C$, we have that ${\uparrow}D\cap{\downarrow}D^{\prime}\neq\emptyset$. We write $V_{g,d}(V_{g}(X))$ for the subset of $V_{r_{g}}(V_{g}(X))$ consisting of the well-directed elements.
@@
Let $\mathcal{A}$ be a Heyting algebra, let $\mathcal{D}$ be a distributive lattice, and let $p:\mathcal{A}\to\mathcal{D}$ be a distributive lattice homomorphism which preserves the relative pseudocomplements from $\mathcal{A}$ (such that $D$ contains them). Let $\mathcal{K}$ be the dual of $V_{GK}(X)$. Then $\mathcal{K}$ is the $\mathsf{KC}$-algebra freely generated by $D$ which preserves the relative pseudocomplements coming from $p$.
@@
Note that the above construction is not guaranteed to provide an extension of the algebra $\mathcal{D}$ – it might be that some elements are identified by the first quotient. However, it is not difficult to show that if $\mathcal{D}$ is a $\mathsf{KC}$-distributive lattice, then it embeds into $\mathcal{K}(\mathcal{D})$ (by showing, for instance, that the composition with the root maps to $\mathcal{D}$ is always surjective). We also note that one could at this point show several of the same facts we did for Heyting algebras for $\mathsf{KC}$-algebras.
@@ 5.2 LC-Algebras
We now consider the case of $\mathsf{LC}$ algebras. As is well-known, such a variety is locally finite. As such it will be interesting to both study the general construction and to outline how this might look when the starting algebras are finite.
First, let us say that a Priestley space $X$ is prelinear if for each $x\in X$, ${\uparrow}x$ is a linear order.
@@
Let $g:X\to Y$ be a Priestley morphism. We say that $C\in V_{g}(X)$ is a linearised closed, rooted and $g$-open subset, if $C$ is a chain. Denote by $V_{g,L}(X)$ the set of linearised closed, rooted and $g$-open subsets.
@@
And we cosntruct the following complex:
@@
Let $Z$ be a prelinear Priestley space, and let $h:Z\to X$ be a $g$-open, continuous and order-preserving map. Then the unique $r_{g}$-open, continuous and order-preserving map $h^{\prime}$ making the obvious triangle commute factors through $V_{g,L}(X)$.
@@
Denote by $V_{g,L}^{\infty}(X)$ the projective limit under the above complex. Then in the same way as before, we can prove:
@@
The following is within the scope of the methods of Ghilardi, using our modifications at hand for the case of $\mathsf{LC}$. We found it in the context of analysing the infinite case, though a simpler finite combinatorial proof is presented here.
Consider a finite poset $P$. Because of local finiteness, we know that $V^{\infty}(P)\cong V^{n}(P)$ must hold for some finite $n$ (otherwise the construction would add infinitely many new points). It is thus natural to wonder whether this construction can stabilise in a single step, i.e., whether after a single application of $V_{L}(X)$ the result will be the final result. Even for very simple posets it becomes apparent that this is not the case. However, surprisingly, two iterations seem to be enough.
@@
Let $C,D\in V_{L}^{2}(X)$, and assume that $V_{L}^{2}(X)$ is prelinear. Suppose that $r(C)=r(D)$ and that $C$ and $D$ are comparable. Then $C=D$.
@@
Assume that $C\nleq D$, without loss of generality. By assumption, then $D<C$. Let $M\in D$ such that $M\notin C$. Hence we may assume that $M\neq r(C)=r(D)$. Consider $T={\uparrow}M\cap D$. Then $D\leq T$. Then there are two cases:
  1. 1.
    If $T\leq C$, then $r(C)\in T$, a contradiction to the fact that $r(C)<M$.
  2. 2.
    If $C\leq T$, then $M\in C$, a contradiction to the assumption.
In both cases we find a contradiction, and hence conclude that $C=D$. ∎
@@
Let $X$ be a finite poset. Assume that $V_{L}^{2}(X)$ is prelinear. Then $V_{L}^{3}(X)\cong V_{L}^{2}(X)$.
@@
Let $X$ be a finite poset. Then $V_{L}^{2}(X)$ is prelinear.
@@
We show, by induction on the cardinality of $x\in V_{L}^{2}(X)$, that ${\uparrow}x$ is linear. Now first, notice that if $x$ has cardinality $1$, then it must be of the form $\{T_{i}\}$ for some $T_{i}\in V_{L}(X)$. Then note that $T_{i}$ must be a singleton: if it had more than one element, say $T_{i}=\{x_{0},...,x_{n}\}$, then $T_{i}\leq\{x_{i},...,x_{n}\}$, for some $x_{i}\neq r(T_{i})$, and this would violate $r$-openness. This means that the elements of cardinality $1$ are exactly the elements of depth $1$. Now assume that $x$ is some element which is of depth $2$, i.e., all of its proper successors are of depth $1$. Then we claim it can have at most one such successor. Otherwise $\{a\}\in x$ and $\{b\}\in x$ for some elements $a,b$. But this is a contradiction, since $x$ is assumed to be linearly ordered, and neither $\{a\}\leq\{b\}$ nor $\{b\}\leq\{a\}$. Hence, since $x$ is of depth $2$, it must contain some unique additional rooted set $T_{i}$, such that whenever $T_{i}\leq M$, then $r(M)=r(T_{i})$ or $r(M)=\{c\}$ for the unique successor of $x$. This easily can be seen to entail that $T_{i}$ must then have exactly two elements. Proceeding in the same way, if $x$ is some element of depth $3$, all of its proper successors are of depth $2$, then it can have at most one such successor by similar arguments, and it must have exactly $3$ elements: one of each cardinality from $1$ to $3$. The argument generalizes straightforwardly for any $n$. But this implies that then ${\uparrow}x$ is linear for any $x$. ∎
Putting these Propositions together, we obtain the following result, which is interesting from several points of view:
@@
Let $X$ be a finite poset. Then $V_{L}^{2}(X)\cong V_{L}^{\infty}(X)$.
@@
The above theorem, coupled with some of the theory we have developed for Heyting algebras, and which could easily be generalised for Gödel algebras, can have some computational uses: given two Gödel algebras, to compute their coproduct, we can look at their dual posets, form the product, and apply $V_{L}^{2}$, to obtain the resulting product. This provides an alternative perspective from the recursive construction discussed in [11, 1]. Moreover, it provides the following interesting property of Gödel logic:

### v3

@@ 4.2 Coproducts and Colimits of Heyting Algebras
are surjective, since for each $x\in X$, $f(x)=g(y)$ for some $y\in Y$, by surjectivity of $g$, and vice-versa. Certainly then the unique lifting from $V_{G}(X\times_{Z}Y,p_{Y},p_{X})$ to the projections will likewise be surjective, and both will be p-morphisms, showing that amalgamation holds. ∎
4.3 Esakia Duality from Priestley Duality
As a final application we show how Esakia duality for Heyting algebras can be presented coalgebraically on the basis of Priestley by using the previous functor. For that purpose, given a Priestley space $X$, recall the map
$\varepsilon:V_{G}(X)\to X,$
the counit of the adjunction, which is simply $\pi_{0}$. Moreover, if $X$ is an Esakia space, recall the map
$\eta:X\to V_{G}(X)$
the unit of the adjunction, which is unique p-morphism lifting the identity. Explicitly, $\eta$ maps $x\in X$ to the tuple
$(x,{\uparrow}x,{\uparrow}({\uparrow}x),...).$
We write $\mu$ for the map $V_{G}(\eta)$. We will now see the functor $V_{G}$ as the composition $V_{G}\circ I$, where $I$ is the inclusion, i.e., we will see it as a comonad on the category of Priestley spaces. In this way, recall that a coalgebra for this comonad is a Priestley space $X$, together with a monotone and continuous function $f:X\to V_{G}(X)$ subject to the following diagrams commuting:
${X}$${V_{G}(X)}$${X}$${V_{G}(X)}$${X}$${V_{G}(X)}$${V_{G}(V_{G}(X))}$$\scriptstyle{f}$$\scriptstyle{1_{X}}$$\scriptstyle{\varepsilon_{X}}$$\scriptstyle{f}$$\scriptstyle{f}$$\scriptstyle{\mu}$$\scriptstyle{V_{G}(f)}$
Figure 8: Coalgebras for the $V_{G}$ comonad
Note that the former diagram implies that $\varepsilon_{X}(f(x))=x$. We will now show the following:
Theorem 37.
Let $X$ be a Priestley space. Then the following are equivalent:
  1. 1.
    $X$ is an Esakia space;
  2. 2.
    $X$ admits a $V_{G}$-coalgebra structure.
  3. 3.
    The map ${\uparrow}:X\to V(X)$ is a monotone and continuous function;
  4. 4.
    The map ${\uparrow}:V(X)\to V(X)$ is a monotone and continuous function.
Proof.
(1) implies (2): if $X$ is an Esakia space, then it admits a free $V_{G}$-coalgebra structure, namely the unit map $\eta:X\to V_{G}(X)$.
Now we show that (2) implies (3). Assume that $f:X\to V_{G}(X)$ is a coalgebra for this comonad. Let $U\subseteq X$ be a clopen subset. We want to show that
${\uparrow}^{-1}[[U]]=\{x:{\uparrow}x\in[U]\}=\Box U$
is clopen. Consider the subset
$U^{*}\coloneqq\{z\in V_{G}(X):z(0)\in U\text{ and }z(1)\in[U]\}.$
Note that this is a clopen subset of $V_{G}(X)$. Hence $f^{-1}[U^{*}]$ is a clopen subset of $X$. Now note that $f(x)=(x,C_{0}^{x},C_{1}^{x},...)$. If $x\leq y$, then $f(x)\leq f(y)$, so since $C_{0}^{x}\leq C_{0}^{y}$, we have that $C_{0}^{y}\subseteq C_{0}^{x}$, and so, $y\in C_{0}^{x}$. In other words, $C_{0}^{x}={\uparrow}x$.
Now then, note that
$f^{-1}[U^{*}]=\Box U.$
Indeed if $x\in\Box U$, then ${\uparrow}x\in[U]$, then by what we just discussed, $f(x)(0)\in U$ and $f(x)(1)\in[U]$, so by the above remark, $f(x)\in U^{*}$. Similarly, if $f(x)\in U^{*}$ it is clear to see that $x\in\Box U$. Hence because $f$ is continuous, we have that $\Box U$ is clopen, as desired.
Now we show that (3) implies (4). Note that this is the map sending a set $C$ to the set ${\uparrow}C$. Indeed if $U$ is clopen in $X$, then
${\uparrow}^{-1}[[U]]=\{C\in V(X):{\uparrow}C\in[U]\}=\{C\in V(X):C\subseteq\Box U\}=[\Box U]$
and we have that $\Box U$ is clopen, since it is exactly ${\uparrow}^{-1}[[U]]$ via the continuous map ${\uparrow}:X\to V(X)$.
Finally, to show (4) implies (1), i.e., that $X$ is an Esakia space, assume that $U$ is clopen. We want to show that ${\downarrow}U$ is open. By pulling $\langle U\rangle$ along ${\uparrow}U$ we have that
${\uparrow}^{-1}[\langle U\rangle]=\{C:{\uparrow}C\cap U\neq\emptyset\}=\{C:C\cap{\downarrow}U\neq\emptyset\}=\langle{\downarrow}U\rangle.$
[…]
$\bigcup_{i\in I}[V_{i}]\cap\langle W_{j_{0},i}\rangle\cap...\cap\langle W_{j_{k},i}\rangle$
for clopens $V_{i}$ and $W_{j,i}$. Now if $x\in{\downarrow}U$, then $\{x\}\in\langle{\downarrow}U\rangle$, so $\{x\}\in[V_{i}]\cap\langle W_{j_{0},i}\rangle\cap...\cap\langle W_{j_{k},i}\rangle$ for some $j,i$; i.e., $V_{i}\cap W_{j_{0},i}\cap...\cap W_{j_{k},i}$ will be an open neighbourhood of $x$ entirely contained in ${\downarrow}U$. This shows that ${\downarrow}U$ is open, showing that $X$ is an Esakia space.∎
As a consequence we obtain that Esakia spaces can be seen, coalgebraically, as arising exactly from Priestley spaces by applying the $V_{G}$ construction. In categorical terms this simply says that the inclusion $I:\mathbf{Esa}\to\mathbf{Pries}$ is a comonadic adjunction. As discussed in my paper with Nick Bezhanishvili [coalgebraicintuitionisticmodal], this perspective can be taken further by looking at the ways in which coalgebraic representations for Priestley spaces can be lifted to coalgebraic representations for Esakia spaces, with applications in the study of modal Heyting algebras of various sorts [^4] In abstract terms, one can use the adjunction $V_{G}\vdash I$ to lift any endofunctor $F:\mathbf{Pries}\to\mathbf{Pries}$ to an endofunctor $\overline{F}:\mathbf{Esa}\to\mathbf{Esa}$, and know that if $F$ is a comonad (e.g., represents a class of coalgebras dual to some appropriate class of modal distributive lattices), then $\overline{F}$ will again be a comonad – indeed, $\overline{F}$ can be computed as the left Kan extension of $F$ along $V_{G}$. I am indebted to Alexander Kurz for this observation..
@@ 5 Subvarieties of Heyting algebras
As is well known, unlike products, coproducts of algebras can vary drastically when one moves from a variety $\mathbf{K}$ to a subvariety $\mathbf{K}^{\prime}$. As such, in this section, we briefly discuss with a few examples how the above construction can be adapted to handle some well-known subvarieties of $\mathbf{HA}$. We will begin by outlining how this works for Boolean algebras, which will allow us to recover the well-known construction of the free Boolean extension of a distributive lattice [gehrkevangoolbooktopologicalduality]. We will then provide analogous free constructions of $\mathsf{KC}$-algebras and $\mathsf{LC}$-algebras [^5] Since completing this paper we have learned of independent work by Carai [caraifreegodelalgebras] outlining the construction of Free Gödel algebras. For a comparison with that approach one can check Section 6 of that paper, which explains the relationship between Carai’s construction and ours.; we leave a detailed analysis of the full scope of the method for future work.
@@
they are also sometimes called WLEM-algebras (For Weak Excluded Middle), DeMorgan Algebras, amongst several other names. They appear importantly in settings like Topos Theory (see e.g. [caramellotopostheoretic]). In turn, $\mathsf{LC}$-algebras are Heyting algebras axiomatised by
@@
they are sometimes called “Gödel algebras” or “Gödel-Dummett algebras” as well as “prelinear Heyting algebras” (the latter is due to the fact that the subdirectly irreducible elements in such a variety are chains). They are of key interest in the study of many-valued logic and substructural logic.
5.1 Boolean algebras
Our general method in handling subvarieties will be to take the Vietoris complex construction, and show that by quotienting by additional equations, one can obtain sub-complexes which limit to the free algebra in the desired variety. To illustrate this, we consider the simple case of Boolean algebras: let $\mathcal{D}$ be a distributive lattice. Consider $\mathbf{F}_{DLat}(\{(a,b):a,b\in D\})/\Theta\cup\{p\vee\neg p=1\}$, the free algebra quotiented as before, with the addition of the equation $p\vee\neg p$, with which one forces Booleannness. Dually, this takes a given subspace of the space $V_{1}(X)$:
Lemma 38.
Let $X$ be a Priestley space. Given a point $C\in V_{1}(X)$, we have that $C\in[\neg U]\cup[U]$ for each $U$ if and only if $C$ is a singleton.
Proof.
If $C$ is a singleton this is clear. Now assume that $C$ has two points $x,y$. Then either $x\nleq y$ or $y\nleq x$; without loss of generality assume the former. Then $x\in U$ and $y\notin U$ for some clopen upset. So $C\notin[U]$ and $C\notin[\neg U]$, as desired. ∎
Using Lemma 38 one can prove the following:
Theorem 39.
For each Priestley space $X$, $B_{1}(X)\cong X$ as Stone spaces, and $B_{1}(X)$ is isomorphic to the free Stone space generated by $X$.
Proof.
The isomorphism is clear; the freeness follows by the same arguments as in Lemma 20. ∎
We thus get that $B_{1}(X)$, the one-step extension from distributive lattices to Boolean algebras, actually coincides with the well-known free Boolean extension, the left adjoint to the inclusion of Boolean algebras in distributive lattices.
It is also worthy to note that Boolean algebras have a peculiar feature as far as our construction $V_{G}$ goes: if $X$ is a discrete Priestley space (definitionally equivalent to a Stone space), then $V_{G}(X)\cong X$, and in fact $V_{1}(X)\cong X$ already (since rooted subsets of singletons have to simply be singletons as well). Stone spaces also exhibit some further interesting properties in this respect:
Theorem 40.
Let $X$ be a Priestley space; then the following are equivalent:
  1. 1.
    $X$ is a Stone space;
  2. 2.
    $V_{1}(X)\cong X$;
  3. 3.
    For some $n$, $V_{n}(X)\cong V_{n+1}(X)$ where the root map witnesses the isomorphism.
  4. 4.
    $V_{G}(X)\cong X$ through the projection $\pi_{0}$, i.e., the counit of the adjunction is an isomorphism for $X$.
Proof.
By the previous remark (1) implies (2), and similar considerations mean that (2) implies (3). We show that (3) implies (1).
Assume that $X$ is not a Stone space, i.e., there is a pair of points $x,y$ such that $x\leq y$. Note that then $x<y$ is a subposet of $X$. Recall from Example 10 that from such a poset one constructs the whole Rieger-Nishimura ladder, and there is throughout each layer of the construction a point $z_{n}$ such that $r$ does not provide an isomorphism. Since all sets involved are finite, this means that $V_{n}(X)$ will always likewise contain such a point $z_{n}$ where $r$ cannot provide an isomorphism.
Next note that certainly (1) implies (4). Now suppose that $X$ is not a Stone space. Then as noted $V_{1}(X)\not\cong X$, so there are at least two elements $C_{0},C_{1}\in V_{1}(X)$ which have the same root. But then we can construct extensions of such elements to elements of $V_{G}(X)$ whilst sharing the same root, i.e., the projection from $V_{G}(X)$ is not injective. ∎
The above theorem does not rule out that there can be other Priestley spaces $X$ such that $V_{G}(X)\cong X$, but it implies that such isomorphisms will not be natural with respect to the adjunction. Such a question is left as an interesting open problem.
@@ 5.2 KC-algebras
Let $(X,\leq)$ be a Priestley space. Let $R=(\leq\cup\leq^{\frown})^{*}$.We say that a subset $C$ is order-connected if for each $x,y\in C$, we have $xRy$. We say that $C$ is an order-component if it is a maximal order-connected subset.
@@
Let us first outline our construction in algebraic terms. Given an appropriate distributive lattice $\mathcal{D}$, consider the sequence:
@@
where we denote the inclusions as $i_{i,j}$. As noted before, the freely generated Heyting algebra is simply the directed limit of this sequence. In order to generate a $\mathsf{KC}$ algebra, we need to quotient this sequence at appropriate steps. Namely, we need to replace $H^{2}(\mathcal{D})$ by the following:
@@
We then construct the alternative sequence:
@@
We will now dually study what this construction looks like.
@@
Let $g:X\to Y$ be a continuous and order-preserving map. Let $C\in V_{r_{g}}(V_{g}(X))$ be an element of the second stage of the $g$-Vietoris complex over $X$. We say that $C$ is well-directed if whenever $D,D^{\prime}\in C$, we have that ${\uparrow}D\cap{\downarrow}D^{\prime}\neq\emptyset$. We write $V_{g,K}(V_{g}(X))$ for the subset of $V_{r_{g}}(V_{g}(X))$ consisting of the well-directed elements.
@@
Let $\mathcal{A}$ be a Heyting algebra, let $\mathcal{D}$ be a distributive lattice, and let $p:\mathcal{A}\to\mathcal{D}$ be a distributive lattice homomorphism which preserves the relative pseudocomplements from $\mathcal{A}$ (such that $\mathcal{D}$ contains them). Let $\mathcal{K}$ be the dual of $V_{GK}(X)$. Then $\mathcal{K}$ is the $\mathsf{KC}$-algebra freely generated by $\mathcal{D}$ which preserves the relative pseudocomplements coming from $p$.
@@
Note that the above construction is not guaranteed to provide an extension of the algebra $\mathcal{D}$ – it might be that some elements are identified by the first quotient. However, it is not difficult to show that if $\mathcal{D}$ is dual to a directed Priestley space, then it embeds into $\mathcal{K}(\mathcal{D})$ (by showing, for instance, that the composition with the root maps to $\mathcal{D}$ is always surjective). We could also run similar constructions as those done for Heyting algebras in the previous sections, with a completely parallel theory.
@@ 5.3 LC-Algebras
We finally consider the case of $\mathsf{LC}$ algebras. As is well-known, such a variety is locally finite, which means that we ought to expect our step-by-step construction to in fact stabilise at a given finite step when the original algebra is finite. As we will see, this is true in fact in greater generality. Let us say that a Priestley space $X$ is prelinear if for each $x\in X$, ${\uparrow}x$ is a linear order.
@@
Let $g:X\to Y$ be a monotone and continuous function. We say that $C\in V_{g}(X)$ is a linearised closed, rooted and $g$-open subset, if $C$ is a chain. Denote by $V_{g,L}(X)$ the set of linearised closed, rooted and $g$-open subsets.
@@
And we cosntruct the following sequence, dual to a Vietoris-like complex:
@@
Let $Z$ be a prelinear Priestley space, and let $h:Z\to X$ be a $g$-open, continuous and order-preserving map. Then the unique $r_{g}$-open, continuous and order-preserving map $h^{\prime}$, making the obvious triangle commute, factors through $V_{g,L}(X)$.
@@
Denote by $V_{G}^{L}(X)$ the projective limit of the sequence
$(X,V_{1}^{L}(X),V_{2}^{L}(X),...)$
where $V_{i}^{L}(X)$ is defined as in Definition 16 except that from the second step onwards we use $V_{g,L}(-)$ rather than $V_{g}(-)$. Then in the same way as before, we can prove:
@@
Consider a finite poset $P$. Because of local finiteness, we know that $V_{G}^{L}(P)\cong V_{n}^{L}(P)$ must hold for some finite $n$ (otherwise the construction would add infinitely many new points). By Theorem 40, $V^{L}_{1}(X)\cong X$ can never hold for such a poset, since $V_{1}^{L}(-)=V_{1}(-)$. But it is reasonable to ask whether $V_{n}^{L}(X)\cong V_{n+1}^{L}(X)$ holds for any poset. Notably, we will see that in fact the construction always stabilises in two steps.
@@
Let $C,D\in V_{L}^{2}(X)$. Suppose that $r(C)=r(D)$ and that $C$ and $D$ are comparable. Then $C=D$.
@@
Assume that $x\in V_{L}^{2}(X)$, and let $R$ be its root. Suppose that $A\in x$ is arbitrary. Then we claim that
$A={\uparrow}r(A)\cap R.$
Indeed, if $b\in A$, then since $R\leq A$, $b\in R$, and by definition $b\in{\uparrow}r(A)$. Conversely, if $b\in R\cap{\uparrow}r(A)$, then $R\leq{\uparrow}b\cap R$ so there is some $B\in x$ such that $r(B)=b$. Note that since $x$ is a chain, we have that $A\leq B$ or $B\leq A$; in the first case $b\in A$, in the latter, since $r(A)\leq b$, we have $r(A)=b$, which again means that $b\in A$. In either case we obtain equality.
Now assume that $C,D\in x$ and $r(C)=r(D)$. Then by what we have just showed:
$C={\uparrow}r(C)\cap R={\uparrow}r(D)\cap R=D.$
@@
Let $X$ be a Priestley space. Assume that $V_{L}^{2}(X)$ is prelinear. Then $V_{L}^{3}(X)\cong V_{L}^{2}(X)$.
@@
Let $X$ be a Priestley space. Then $V_{L}^{2}(X)$ is prelinear.
@@
Assume that $x\leq y$ and $x\leq z$, and look at $r(y)$ and $r(z)$. Since these are in $x$, which is a chain, assume that $r(y)\leq r(z)$. Then we claim that $y\leq z$. Indeed, assume that $A\in z$. Then $r(z)\leq A$, so $r(y)\leq A$. Because $y$ is $r$-open, let $A^{\prime}$ be such that $A^{\prime}\in y$, and $r(A)=r(A^{\prime})$. Then $A^{\prime}\in x$, and since $A\in z$, $A\in x$ as well; so by Lemma 51, we have that $A=A^{\prime}$. This shows that $A\in y$, which shows the result. ∎
Putting these Propositions together, we obtain the following result:
@@
Let $X$ be a Priestley space. Then $V_{L}^{2}(X)\cong V_{L}^{\infty}(X)$.
@@
The above theorem, coupled with some of the theory we have developed for Heyting algebras, and which could easily be generalised for Gödel algebras, can have some computational uses: given two Gödel algebras, to compute their coproduct, we can look at their dual posets, form the product, and apply $V_{L}^{2}$, to obtain the resulting product. This provides an alternative perspective from the recursive construction discussed in [DAntona2006, aguzzoligodelcoproducts], and can be compared with the work of Carai in [caraifreegodelalgebras]. In logical terms, it implies the following interesting property of Gödel logic: